Fermi National Laboratory
 

Accelerator Update for February 26, 2007 - February 28, 2007

 

Monday February 26

The day shift began with Operations monitoring Tevatron (TeV) store 5245 (79.1E30) and the Recycler stash (208.3E10), with the Antiproton Source (Pbar) stacking (44.2E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

At 10:47 AM, the NuMI beam permit tripped off due to a quadrupole’s (E:QF112) current drifting low. It reset okay.

Switchyard (SY) experts began a study period at 11:37 AM.

A TeV expert began a Tevatron Electron Lens (TEL1) study at 2:07 PM.

Operations reported at 4:14 PM that the Recycler had stashed antiprotons four times during the day shift.

At 4:54 PM, the NuMI beam permit tripped off due to a quadrupole’s (E:QD108) current drifting low. It reset okay.

The TEL1 study ended at 5:44 PM.

Operations terminated store 5245 at 9:52 PM.

Operations put the TeV into shot setup at 10:07 PM.

At 11:27 PM, Operations reported that the shot setup had been put on hold for 30 minutes because the antiprotons were not cooling properly. Operators dumped the protons that had already been loaded into the TeV.

Operations reported at 11:59 PM that the Recycler had stashed twice during the evening shift.

Tuesday February 27

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (292.7E10) and stack (37.6E10), and with NuMI taking beam.

Operations established store 5246 at 12:55 AM with an initial luminosity of 262.27E30.

Operations reported at 7:59 AM that the Recycler had stashed three times during the midnight shift.

At 12:28 PM, NuMI tripped off for the first of many times in two hours. It reset.

Operations reported at 4:05 PM that the Recycler had stashed three times during the day shift.

At 4:23 PM, a Linac RF station (LRF4) tripped off for the first of four times in six minutes. Linac experts investigated. They reported that the power amplifier was weak and would last for more than a few days.

At 5:10 PM, BRF6 tripped off. It reset.

At 5:53 PM, NuMI tripped off for the first of three times in nine minutes. It reset.

Operations reported at 11:59 PM that the Recycler had stashed three times during the evening shift.

Wednesday February 28

The midnight shift began with Operations monitoring store 5246 (41.25E30) and the stash (339.6E10), with Pbar stacking (54.6E10), and with NuMI taking beam.

Operations terminated store 5246 at 2 AM.

Operations established a dry squeeze at 6:40 AM.

At 6:48 AM, the NuMI horn breaker tripped off. It reset, but then began tripping off on stripline ground faults until its breaker tripped off again. Operations contacted experts.

Per NuMI expert instructions, Operations put NuMI in the horn off running mode at 7:08 AM.

Operations reported at 7:56 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Wednesday

The plans for today are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/01/2007    email Fermilab